freon:

MATHEMATICS At the ancient Athens Olympic games, in a duel between two runners, Portheus and Morpheus, they were made to start running in opposite direction from diametrically opposite ends of a circular race track of length (circumference) 2 kms. The first time they met was after 24 minutes. If the distance between them exactly 'n' minutes after they start is equal to a quarter of the length of the track, which of the following is a possible value of 'n'?

abound:

Why is it titled mathematics but is under writing

freon:

i'm not getting option for subject

1 attachment
abound:

n word

freon:

nice?

abound:

@freon wrote:
nice?
nice....

abound:

wdym nice

notmeta:

want the explanation with answer or without?

freon:

itll be much

@notmeta wrote:
want the explanation with answer or without?
helpful with the explaination!

Luigi0210:

Abound stop spamming the post. And meta, behave..

notmeta:

@luigi0210 wrote:
Abound stop spamming the post. And meta, behave..
I am just trying to answer :)

notmeta:

To find a possible value of 'n', we need to analyze the motion of two runners, Portheus and Morpheus, on a circular race track. By using the information that they meet after 24 minutes and the distance between them after 'n' minutes is a quarter of the track length, we can set up equations based on their speeds and distances to determine the possible value of 'n'. The equation we can use to solve for 'n' is: (x + y) * 24 = 2 (for when they meet after 24 minutes) (x + y) * n = 0.5 (for when the distance between them after 'n' minutes is a quarter of the track length)

freon:

@notmeta wrote:
To find a possible value of 'n', we need to analyze the motion of two runners, Portheus and Morpheus, on a circular race track. By using the information that they meet after 24 minutes and the distance between them after 'n' minutes is a quarter of the track length, we can set up equations based on their speeds and distances to determine the possible value of 'n'. The equation we can use to solve for 'n' is: (x + y) * 24 = 2 (for when they meet after 24 minutes) (x + y) * n = 0.5 (for when the distance between them after 'n' minutes is a quarter of the track length)
shouldn't first equation has 1 as rhs as diametrically opp means half of circumference?

freon:

got the answer from your explaination! tysm!

@notmeta wrote:
To find a possible value of 'n', we need to analyze the motion of two runners, Portheus and Morpheus, on a circular race track. By using the information that they meet after 24 minutes and the distance between them after 'n' minutes is a quarter of the track length, we can set up equations based on their speeds and distances to determine the possible value of 'n'. The equation we can use to solve for 'n' is: (x + y) * 24 = 2 (for when they meet after 24 minutes) (x + y) * n = 0.5 (for when the distance between them after 'n' minutes is a quarter of the track length)
got the answer from your explaination! tysm!